OpenStudy (lena772):

Winnie poured 14 cups of water into a rectangular container measuring 13 inches by 7 inches by 6 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points)

ganeshie8 (ganeshie8):

start by converting 6 centimeters to inches

OpenStudy (lena772):

2.54 cm in an inch right?

ganeshie8 (ganeshie8):

\(\large 6 ~cm = 6~ cm\times \frac{1~ in}{2.54~cm} = 6\times \frac{1~ in}{2.54} = 2.36 ~in \)

OpenStudy (lena772):

:)

ganeshie8 (ganeshie8):

so the dimensions of container are :- length = 13 in width = 7 in height = 2.36 in

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

for part A, use the formula for Volume of regular prism

OpenStudy (lena772):

214.76 in^2

OpenStudy (lena772):

l*w*h right?

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

thats the max volume of water or ANYTHING the container can hold

ganeshie8 (ganeshie8):

btw, volume is measured in cubics... 214.76 in^3

ganeshie8 (ganeshie8):

you must never say volume is 214.76 in^2, its always in^3 ok

ganeshie8 (ganeshie8):

in^3 cm^3 ft^3 m^3 ....

OpenStudy (lena772):

ok

ganeshie8 (ganeshie8):

wat about part B

ganeshie8 (ganeshie8):

Winnie poured 14 cups of water into a rectangular container [1 cup = 14.44 cubic inches;

ganeshie8 (ganeshie8):

So total how many cubic inches water was poured into the container ?

OpenStudy (lena772):

202.16in^3

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

for part C :- say \(h\) is the height of water in the container wid length and width as : 13 and 7

ganeshie8 (ganeshie8):

then, Volume of water = 13*7*h

ganeshie8 (ganeshie8):

you knw volume of water from Part B, use it

ganeshie8 (ganeshie8):

202.16 = 13*7*h

ganeshie8 (ganeshie8):

you can find h

OpenStudy (lena772):

202.16/(13*7)=h

OpenStudy (lena772):

2.22=h

OpenStudy (lena772):

Is that right?

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\) 2.22 in